OpenStudy (anonymous):

find the function f(x) if f'(x)=2x-3 and f(1)=10

OpenStudy (anonymous):

If this involves integrals: $$\int {2x-3} dx = x^2 - 3x$$

OpenStudy (anonymous):

?

OpenStudy (anonymous):

f(x) = x^2 - 3x +12 f'(x) = 2x - 3 f(1) = 10